OpenStudy (anonymous):

Can someone please help? A wooden pyramid, 9 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 4 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids.

OpenStudy (cwrw238):

height of new pyramid = 4*9 = 36 ins the ratio of the heights is 4 : 1 - the ratio of the volumes is 4^3 : 1^3 = 64: 1

OpenStudy (anonymous):

Thank you! :)

OpenStudy (cwrw238):

yw